### v2

@@ 4 Simulation studies: American put options
where the stock price follows a one-dimensional geometric Brownian motion: $\operatorname{d}\!X_{t}=\rho X_{t}\operatorname{d}\!t+\sigma X_{t}\operatorname{d}\!W_{t}$ with $\rho$ being the risk-free rate and $\sigma$ being the volatility of the stock price, and the (time-invariant) payoff function is $g(x)=\left(\Gamma-x\right)^{+}$ with $\Gamma$ being the strike price. In the RL setting, only the volatility $\sigma$ is unknown to the agent.
Due to the presence of the exponential time discounting in the objective function, (3.4) and (3.7) in Algorithm 1 and 2 need to be respectively modified to (refer to Section 5.1 in Jia and Zhou, 2022a, )
@@
Number of training steps $N$ 1000 or 5000
@@ 4.2 Learned value functions
With the choice of the parameter values in Table 1, we implement training for 1000 steps by the offline ML Algorithm 1 and for 5000 steps by the online TD(0) Algorithm 2. These are carried out within the PyTorch framework. Each algorithm spends about 0.5s for one step training on a Windows 11 PC with an Intel Core i7-12700 2.10 GHz CPU and 32GB RAM by CPU computing only.
Figure 1 shows the learning results by Algorithms 1 and 2 regarding the exploratory value functions at six selected decision time points $t\in\{t_{l}:l=0,10,20,30,49\}=\{0,0.2,0.4,0.6,0.8,0.98\}$. In each panel, we plot five lines with the stock price ranging from zero to twice the initial stock price. The solid red line (with label “Option”) shows the theoretical (omniscient) values of $v(t,x)-g(x)$ derived by finite difference. The dotted solid blue, orange and green lines (with label “Dong (2024)”, “ML” and “TD(0)”) are respectively the outputs $v-g$ from the NN by the TD error algorithm in Dong, (2024) and by the ML and TD(0) algorithms in this paper. The dotted yellow vertical line (with label “$S^{*}$”) refers to the specific stock price that separates the stopping and continuation regions at each time, which is also derived by finite difference.
Figure 1: Learning results of $v-g$ under the TD algorithm in Dong, (2024), the offline ML algorithm and the online TD(0) algorithm in this paper at six selected decision time points: $\{0,10,20,30,40,49\}$-th discretized time points. The solid red lines are the theoretical values of $v-g$ and the dotted yellow vertical lines specify the stock price on the theoretical free boundary, both derived by finite difference. The training parameters of Dong (2024) are specified in Section 4.6.
@@ 4.3 Execution policy
Next, we calculate the accuracy rate of classification at each time point as
#{correctly classified sample $\frac{\#\left\{\text{correctly classified sample points}\right\}}{M_{\mathrm{test}}}.$
It turns out that both algorithms achieve accuracy rates of higher than $95\%$ for all discretized time points. [^11] One never exercises an American put option when the stock price is greater than the strike price; so actually we only need to study the points below the black dotted lines in Figure 2. With this in mind the accuracy rates achieved are even higher.
@@ 4.4 Option price
The relative error is defined as $\frac{|\mathcal{P}-5.317|}{5.317}$. We calculate the relative error on the testing dataset every 10 training steps. The learning curves of option price from the two methods described above are plotted in Figure 3.
@@ 4.5.1 Penalty factor
We examine the effect of penalty factor on the learned option price in this subsection. First, it follows from (3.6) that $500<K\leqslant\frac{1}{\Delta t}=$ 50 in our setting. We pick $K\in\left\{1,5,10,20,30,40,50\right\}$ and keep the other model parameter values unchanged as in Table 1. Figure 4 shows the learning curves of option prices with a combination of the two algorithms and the two alternative ways. As expected, the higher the value of $K$, the less the relative error achieved by both algorithms. However, when $K$ increases to a certain level, a further increase no longer helps decrease the error, in which case the sampling error dominates.
@@ 4.5.2 Temperature parameter
The temperature parameter controls the weight on exploration, and thus is important for learning. Figure 5 shows the learning curves of option prices when $\lambda\in\left\{0.1,0.5,1,3,5,10\right\}$, while keeping the other model parameter values as in Table 1.121212In Appendix C, we also present the results for a higher penalty factor $K=50$. We observe that in general, the smaller the temperature parameter, the more accurate but the lower the learning speed for convergence, which is especially evident in the TD(0) case. [^13] Similar results are also derived in Dong, (2024). An implication of this observation is that if the computing resource is limited so the number of training steps is small, we can increase $\lambda$ to accelerate convergence. Intriguingly, in Figure 5, the errors are not monotonic in $\lambda$ when $\lambda\leqslant 1$ in the two cases using $\mathcal{P}_{\mathrm{stopping}}$. We conjecture that it is because the effect of $\lambda$ is dominated by the sampling error when $\lambda$ is sufficiently small. However, overall, the differences are around $0.3\%$ for $\lambda\in\{0.1,0.5,1\}$, which are sufficiently small.
@@ 4.6 Comparison with Dong (2024)
Both Dong, (2024) and this paper study RL for optimal stopping, but there are major differences as discussed in the introduction section. In particular, there is a considerable technical difference in terms of the entropy terms introduced in the respective RL objectives. It turns out that, when formulated in our setting with the penalty factor $K$, the decision variable $\tilde{\pi}$ in Dong, (2024) is the product of $K$ and the probability of stopping in our paper, i.e., $\tilde{\pi}=K\pi$. The entropy used in Dong, (2024) is $R(\tilde{\pi}):=\tilde{\pi}-\tilde{\pi}\log\tilde{\pi}=R(K\pi)=K\pi-K\pi\log(K\pi),$ whereas we take the differential entropy $-\mathscr{H}(\pi)=-\pi\log\pi-(1-\pi)\log(1-\pi).$ Dong, (2024) explains that $R(\tilde{\pi})$ is proposed by assuming $\tilde{\pi}\Delta t$ to be the probability of stopping in a $\Delta t$-length time interval and then integrating the differential entropy $-\mathscr{H}(\tilde{\pi}\Delta t)=-\tilde{\pi}\Delta t\log\Delta t+\left(\tilde{\pi}-\tilde{\pi}\log\tilde{\pi}\right)\Delta t+o(\Delta t)$ over the whole time horizon while dropping the first divergence term. In other words, it is not the exact differential entropy of the randomized strategy as in our paper.
Next, we compare the performances of the algorithms between the two papers. We use $\mathcal{P}_{\mathrm{stopping}}$ defined in (4.5) for computing option price, as it is the one taken by Dong, (2024). For a fair comparison, we use the TD error employed in Dong, (2024) as the criterion for performance and we learn the option price instead of the early exercise premium. We implement training for 1000 steps for both algorithms and repeat the training for 10 times. Because the scales of the entropy terms used in the two papers are considerably different, we train the two algorithms for different combinations of temperature parameter $\lambda\in\left\{10,1,0.1,0.01,0.001,0.0001\right\}$ and learning rate $\alpha\in\left\{0.1,0.01,0.001,0.0001\right\}$. The other parameters are kept the same as in Table 1 and the simulated stock price samples are the same for both algorithms. We then check the learning speeds of the two algorithms. Since the theoretical option price is 5.317, we examine the learning speeds for all the cases of $(\lambda,\alpha)$ in which the learned option prices are at least 5.29 for a meaningful comparison. We find that our algorithm is always faster than Dong, (2024)’s. Figure 6 is an illustration with $(\lambda,\alpha)=(0.001,0.01)$ (these are the model parameters corresponding to the solid blue line in Figure 1) in Dong’s algorithm and $(\lambda,\alpha)=(1,0.1)$ in our algorithm. The figure shows that both algorithms achieve very low relative errors ($0.39\%$ for our method and $0.44\%$ for Dong’s method) after 1000 training steps, but our algorithm is around 200 steps faster in convergence. One of the reasons may be that compared to the differential entropy in our algorithm, the scale of the entropy term in Dong, (2024) is larger, causing more biases in learning the objective stopping value.
@@
The simulation study reported so far is for an American option, whose price is evaluated under the risk-neutral probability. However, the underlying stock price is in general not observable in the risk-neural world; hence pricing American options is not the best example for illustrating a data-driven approach such as RL. Here, we present another application to demonstrate the effectiveness of our method in which the physical probability measure instead of the risk-neutral one is used (and thus all the relevant data are observable). Specifically, Dai and Dong, (2024) consider the problem of learning an optimal investment policy to maximize a portfolio’s terminal log return in the presence of proportional transaction costs, and turn it into an equivalent Dynkin game which involves optimal stopping decisions of two players. The latter is naturally formulated under the physical measure, and suitable for RL study under the frameworks of both Dong, (2024) and this paper; refer to Appendix D for details. Dai and Dong, (2024) extend the RL algorithm of Dong, (2024) to learn an optimal investment policiy with transaction costs. With the same model parameters as in Dai and Dong, (2024) and a penalty factor $K=5$ (which is a proper choice when the time horizon is equally divided into 10 time intervals and there exist two stopping times in the Dynkin game), we learn how to trade in the Black–Scholes market under the penalty formulation, and compare the learning results of free boundaries and average log return when applying the learned strategies in a 20-year horizon. Similar to the case of the American put option, we choose temperature parameter $\lambda\in\left\{10^{-2},10^{-4},10^{-6}\right\}$ and learning rate $\alpha\in\left\{0.1,0.05,0.01,0.005\right\}$, and keep the other training parameters and structures of neural networks the same as in Dai and Dong, (2024), with 15000 training steps for both algorithms. As noted, in this example the stock price samples are simulated under the physical measure. Our experiments show that among the 12 combinations of $\lambda$ and $\alpha$, Dai and Dong, (2024)’s algorithm achieves higher average log return in the last training step than our algorithm for 2 times, while we achieve better results for 10 times. For comparing the learning speeds, note that the omniscient level of average log return in the 20-year horizon under the testing dataset is around 3.00; so we choose the cases of $(\lambda,\alpha)$ in which the two algorithms achieve average log return values of greater than 2.95. We find that $(\lambda,\alpha)=(10^{-6},0.05)$ and $(\lambda,\alpha)=(10^{-2},0.01)$ result in the fastest convergence of average log return for Dai and Dong, (2024)’s algorithm and ours, respectively. Figure 7 plots the corresponding learning curves of the initial free boundaries and average log returns. Figure 7(a) shows that the convergence of the sell boundary is almost identical for the two algorithms, while the convergence of the buy boundary is faster under our algorithm. Figure 7(b) indicates that the convergence of average log return is around 5000 steps faster under our algorithm. It is interesting to note that the trivial buy-and-hold strategy leads to an average log return value of 2.5; so both the RL algorithms indeed help the agent “beat the market”.

### v3

@@ 4 Simulation studies: American put options
where the stock price follows a one-dimensional geometric Brownian motion: $\operatorname{d}\!X_{t}=\rho X_{t}\operatorname{d}\!t+\sigma X_{t}\operatorname{d}\!W_{t}$ with $\rho$ being the risk-free rate and $\sigma$ being the volatility of the stock price, and the (time-invariant) payoff function is $g(x)=\left(\Gamma-x\right)^{+}$ with $\Gamma$ being the strike price. In the RL setting, the volatility $\sigma$ is unknown to the agent.
Due to the presence of the exponential time discounting in the objective function, (3.3) and (3.6) in Algorithms 1 and 2 need to be respectively modified to (refer to Section 5.1 in Jia and Zhou, 2022a, ):
@@
Number of training steps NN 1000 or 5000
@@ 4.2 Learned value functions
With the parameter values given in Table 1, we implement training for 1000 steps by the offline ML Algorithm 1 and for 5000 steps by the online TD(0) Algorithm 2. These are carried out within the PyTorch framework. Each algorithm spends about 0.5s for one step training on a Windows 11 PC with an Intel Core i7-12700 2.10 GHz CPU and 32GB RAM by CPU computing only.
Figure 1 shows the learning results by Algorithms 1 and 2 regarding the exploratory value functions at six selected decision time points $t\in\{t_{l}:l=0,10,20,30,49\}=\{0,0.2,0.4,0.6,0.8,0.98\}$. In each panel, we plot five lines with the stock price ranging from zero to twice the initial stock price. The solid red line (with label “Option”) shows the theoretical (omniscient) values of v(t,x)g(x)v(t,x)-g(x) derived by finite difference. The dotted solid blue, orange and green lines (with label “Dong (2024)”, “ML” and “TD(0)”) are respectively the outputs vgv-g from the NN by the TD error algorithm in Dong, (2024) and by the ML and TD(0) algorithms in this paper. The dotted yellow vertical line (with label $S^{*}$”) refers to the specific stock price that separates the stopping and continuation regions at each time, which is also derived by finite difference.
Figure 1: Learning results of vgv-g under the TD algorithm in Dong, (2024), the offline ML algorithm and the online TD(0) algorithm in this paper at six selected decision time points: $\{0,10,20,30,40,49\}$-th discretized time points. The solid red lines are the theoretical values of vgv-g and the dotted yellow vertical lines specify the stock price on the theoretical free boundary, both derived by finite difference. The training parameters of Dong (2024) are specified in Section 4.6.
@@ 4.3 Execution policy
Next, we calculate the accuracy rate of classification at each time point as    $\frac{\#\left\{\text{correctly classified sample points}\right\}}{M_{\mathrm{test}}}.$ It turns out that both algorithms achieve accuracy rates of higher than $95\%$ for all discretized time points. [^13] One never exercises an American put option when the stock price is greater than the strike price; so actually we only need to study the points below the black dotted lines in Figure 2. With this in mind the accuracy rates achieved are even higher.
@@ 4.4 Option price
Recalling that the omniscient option price is 5.317, we define the relative error of a learned price $\mathcal{P}$ to be $\frac{|\mathcal{P}-5.317|}{5.317}$. We calculate the relative error on the testing dataset every 10 training steps. The learning curves of option price from the two methods described above are plotted in Figure 3.
@@ 4.5.1 Penalty factor
We examine the effect of penalty factor on the learned option price in this subsection. First, from a numerical implementation perspective, it follows from (3.5) that $0<K\leqslant\frac{1}{\Delta t}=50$ in our setting, since the process $R_{t}$ cannot become negative. We pick $K\in\left\{1,5,10,20,30,40,50\right\}$ and keep the other model parameter values unchanged as in Table 1. Figure 4 shows the learning curves of option prices with a combination of the two algorithms and the two alternative ways. As expected, the higher the value of $K$, the less the relative error achieved by both algorithms. However, when $K$ increases to a certain level, a further increase no longer helps decrease the error, in which case the sampling error dominates.
@@ 4.5.2 Temperature parameter
The temperature parameter controls the weight on exploration, and thus is important for learning. Figure 5 shows the learning curves of option prices when $\lambda\in\left\{0.1,0.5,1,3,5,10\right\}$, while keeping the other model parameter values as in Table 1.141414In Appendix C, we also present the results for a higher penalty factor $K=50$. We observe that in general, the smaller the temperature parameter, the more accurate but the lower the learning speed for convergence, which is especially evident with the TD(0) algorithm. [^15] Similar results are also noted in Dong, (2024). An implication of this observation is that if the computing resource is limited so the number of training steps is small, we can increase $\lambda$ to accelerate convergence. Intriguingly, in Figure 5, the errors are not monotonic in $\lambda$ when $\lambda\leqslant 1$ in the two cases using $\mathcal{P}_{\mathrm{stopping}}$. We conjecture that it is because the effect of $\lambda$ is dominated by the sampling error when $\lambda$ is sufficiently small. However, overall, the differences are around $0.3\%$ for $\lambda\in\{0.1,0.5,1\}$, which are sufficiently small.
@@ 4.6 Comparison with Dong (2024)
[…]
Next, we compare the performances of the algorithms between the two papers. We use $\mathcal{P}_{\mathrm{stopping}}$ defined in (4.5) for computing option price, as it is the one taken by Dong, (2024). For a fair comparison, we use the TD error employed in Dong, (2024) as the criterion for performance and we learn the option price instead of the early exercise premium. We implement training for 1000 steps for both algorithms and repeat the training for 10 times. Because the scales of the entropy terms used in the two papers are considerably different, we train the two algorithms for different combinations of temperature parameter $\lambda\in\left\{10,1,0.1,0.01,0.001,0.0001\right\}$ and learning rate $\alpha\in\left\{0.1,0.01,0.001,0.0001\right\}$. The other parameters are kept the same as in Table 1 and the simulated stock price samples are the same for both algorithms. We then check the learning speeds of the two algorithms. Since the theoretical option price is 5.317, we examine the learning speeds for all the cases of $(\lambda,\alpha)$ in which the learned option prices are at least 5.29 for a meaningful comparison. We find that our algorithm is always faster than Dong, (2024)’s. Figure 6 is an illustration with $(\lambda,\alpha)=(0.001,0.01)$ (these are the model parameters corresponding to the solid blue line in Figure 1) in Dong’s algorithm and $(\lambda,\alpha)=(1,0.1)$ in our algorithm. The figure shows that both algorithms achieve very low relative errors ($0.39\%$ for our method and $0.44\%$ for Dong’s method) after 1000 training steps, but our algorithm is around 200 steps faster in convergence.
@@
The simulation study reported so far is for an American option, whose price is evaluated under the risk-neutral probability. However, the underlying stock price is in general not observable in the risk-neural world; hence pricing American options is not the best example for illustrating a data-driven approach such as RL. Here, we present another application to demonstrate the effectiveness of our method in which the physical probability measure instead of the risk-neutral one is used (and thus all the relevant data are observable). Specifically, Dai and Dong, (2024) consider the problem of learning an optimal investment policy to maximize a portfolio’s terminal log return in the presence of proportional transaction costs, and turn it into an equivalent Dynkin game which involves optimal stopping decisions of two players. The latter is naturally formulated under the physical measure, and suitable for RL study under the frameworks of both Dong, (2024) and this paper; refer to Appendix D for details. Dai and Dong, (2024) extend the RL algorithm of Dong, (2024) to learn an optimal investment policiy with transaction costs. With the same model parameters as in Dai and Dong, (2024) and a penalty factor $K=5$ — which is an appropriate choice as specified in Section 4.5.1 when the time horizon $\left[0,T\right]$ with $T=1$ is equally divided into 10 time intervals (so that $\Delta t=1/10$, and there are two stopping times in the Dynkin game) — we learn how to trade in the Black–Scholes market under the penalty formulation. We then compare the learning results for the free boundaries and the average log return when applying the learned strategies over a 20-year horizon. Similar to the case of the American put option, we choose temperature parameter $\lambda\in\left\{10^{-2},10^{-4},10^{-6}\right\}$ and learning rate $\alpha\in\left\{0.1,0.05,0.01,0.005\right\}$, and keep the other training parameters and structures of neural networks the same as in Dai and Dong, (2024), with 15000 training steps for both algorithms. As noted, in this example the stock price samples are simulated under the physical measure. Our experiments show that among the 12 combinations of $\lambda$ and $\alpha$, Dai and Dong, (2024)’s algorithm achieves higher average log return in the last training step than our algorithm for 2 times, while we achieve better results for 10 times. For comparing the learning speeds, note that the omniscient level of average log return in the 20-year horizon under the testing dataset is around 3.00; so we choose the cases of $(\lambda,\alpha)$ in which the two algorithms achieve average log return values of greater than 2.95. We find that $(\lambda,\alpha)=(10^{-6},0.05)$ and $(\lambda,\alpha)=(10^{-2},0.01)$ result in the fastest convergence of average log return for Dai and Dong, (2024)’s algorithm and ours, respectively. Figure 7 plots the corresponding learning curves of the initial free boundaries and average log returns. Figure 7(a) shows that the convergence of the sell boundary is almost identical for the two algorithms, while the convergence of the buy boundary is faster under our algorithm. Figure 7(b) indicates that the convergence of average log return is around 5000 steps faster under our algorithm. It is interesting to note that the trivial buy-and-hold strategy leads to an average log return value of 2.5; so both the RL algorithms indeed help the agent “beat the market”.
@@
[…]
In addition, in Dong, (2024)’s algorithm, the use of the unnormalized negentropy yields an optimal intensity $\tilde{\pi}=\exp\left(-\frac{V_{\theta}-g}{\lambda}\right)$, which takes values in $0,\infty)$. However, for a given time step size $\Delta t$, $\tilde{\pi}$ must be truncated at $\frac{1}{\Delta t}$ if $\exp\left(-\frac{V_{\theta}-g}{\lambda}\right)>\frac{1}{\Delta t}$, since $\tilde{\pi}\Delta t$ is interpreted as the probability of stopping in his formulation. This situation arises when the $\theta$-parameterized neural network generates a negative $V_{\theta}-g$ and the temperature parameter is small, such as $\lambda\leqslant 10^{-3}$ in the numerical examples. As a result, the loss function minimized by stochastic gradient descent is not the exact TD error in his formulation (see equation (4.4) in Dong, (2024)), introducing a bias into the learning process. By contrast, our algorithm employs differential entropy resulting in an optimal probability $\pi\in[0,1]$, and the loss function minimized by stochastic gradient descent corresponds exactly to the TD error in our formulation. This may be another reason why our algorithm exhibits faster convergence.
5 Simulation studies: High-dimensional examples
Other than the model-free approach that mitigates the model estimation errors, another significant advantage of RL is its ability to scale to high-dimensional problems. We demonstrate this with two examples in this section.
5.1 American geometric average put-type option
The following example is adapted from Section 4.3.2.1 in Becker et al., (2021) and concerns the pricing of an American geometric average put-type option on up to 200 distinguishable stocks in a Black–Scholes market.
As in Becker et al., (2021), we set the dimension (the number of stocks) $d\in\{40,80,120,\allowbreak 160,200\}$, the initial stock price $x_{0}=100^{1/\sqrt{d}}$ for all the stocks, the time horizon $T=1$, the strike price $\Gamma=95$, the risk-free rate $\rho=0.6$, the volatility $\sigma_{i}=\min\{0.04[(i-1)\bmod 40],\,1.6-0.04[(i-1)\bmod 40]\}$ and dividend yield $\delta_{i}=\rho-\frac{\sum_{i=1}^{d}\sigma_{i}^{2}}{d^{2}}(i-\frac{1}{2})-\frac{1}{5\sqrt{d}}$ for each $i\in\left\{1,2,\ldots,d\right\}$. Denote by $X_{t}^{(k)}$ the price of the $k$-th stock at time $t$. The optimal stopping problem is to
$\max_{\tau\in[0,T]}\,\mathbb{E}\left[e^{-\rho\tau}\left(\Gamma-\left[\prod_{k=1}^{d}|X_{\tau}^{(k)}|^{1/\sqrt{d}}\right]\right)^{+}\right].$
In our simulation studies, we set the penalty factor $K$ to be 100 (which corresponds to the time mesh grid $\Delta t=1/100$ as discussed in Section 4.5.1), the temperature parameter $\lambda=1$, the training batch size to be $2^{10}$, and the testing batch size to be $2^{17}$. The initial learning rate is set to be 0.05 and is decreased by a factor 1/10 every 300 training steps. We train 1000 steps under the offline ML Algorithm 1. In Table 2, we show the learning results for different numbers $d$ of underlying stocks. We test the learned price every 10 training steps; so we have 100 observations in total for each $d$. The final observation is listed in the second column of Table 2.
It is worth noting that pricing this high-dimensional American geometric average put-type option is equivalent to pricing a one-dimensional American put option that is independent of the dimension. [^16] Specifically, for the equivalent one-dimensional American put-type option problem, the stock price process in a Black–Scholes market setting has an initial price of 100, a risk-free rate of $\rho$, a dividend yield of $\rho-\frac{\sum_{i=1}^{d}\sigma_{i}^{2}}{2d}-\frac{1}{5}$, and a volatility of $\frac{1}{\sqrt{d}}\sqrt{\sum_{i=1}^{d}\sigma_{i}^{2}}$ under the particular choices of $\sigma_{i}$ and $\delta_{i}$ in the example. Notably, $\frac{\sum_{i=1}^{d}\sigma_{i}^{2}}{d}$ remains constant when the dimension is a multiple of 40; see Section 4.3.2.1 and Proposition 4.3 in Becker et al., (2021). Hence we can easily obtain the corresponding omniscient values: The theoretical price of its European counterpart is 3.565, and the benchmark value for the price of this one-dimensional American put option calculated using the binomial tree method on Smirnov’s website (https://www.math.columbia.edu/~smirnov/options13.html) with 20000 nodes is 6.545. It turns out that our RL algorithm achieves high early exercise premium and is effective for pricing the high-dimensional option with comparable high accuracy compared to the benchmarks in Becker et al., (2021). Moreover, our algorithm is scalable, as Table 2 shows a sublinear growth in runtime with respect to the dimension.
Dimension $d$ Last learned price by our method In Becker et al., (2021) Runtime in sec.
40 6.512 6.512 1838
80 6.523 6.509 2956
120 6.504 6.507 4094
160 6.507 6.504 6042
200 6.493 6.501 7394
Table 2: Comparison of learned prices and runtime for different values of dimension $d$. The initial stock price $x_{0}=100^{1/\sqrt{d}}$, the time horizon $T=1$, the strike price $\Gamma=95$, the risk-free rate $\rho=0.6$, the volatility $\sigma_{i}=\min\{0.04[(i-1)\bmod 40],\,1.6-0.04[(i-1)\bmod 40]$ and dividend yield $\delta_{i}=\rho-\frac{\sum_{i=1}^{d}\sigma_{i}^{2}}{d^{2}}(i-\frac{1}{2})-\frac{1}{5\sqrt{d}}$ for each $i\in\left\{1,2,\ldots,d\right\}$. The payoff function when exercising the option is $g(x)=\left(\Gamma-\left[\prod_{k=1}^{d}|x^{(k)}|^{1/\sqrt{d}}\right]\right)^{+}$. The benchmark value for the price of this option is 6.545.
In the above experiment, the risk-free rate $\rho=0.6$ taken from Becker et al., (2021) is unrealistically large. We now select a more reasonable value and adjust other parameters so that the equivalent one-dimensional American put option remains independent of $d\in\left\{40,80,120,160,200\right\}$. Specifically, we set the risk-free rate $\rho=0.06$, the volatility $\sigma_{i}=\frac{1}{\sqrt{10}}\min\left\{0.04[(i-1)\bmod 40],\,1.6-0.04[(i-1)\bmod 40]\right\}$ and dividend yield $\delta_{i}=\rho-\frac{\sum_{i=1}^{d}\sigma_{i}^{2}}{d^{2}}(i-\frac{1}{2})-\frac{1}{50\sqrt{d}}$ for each $i\in\left\{1,2,\ldots,d\right\}$. With these modifications, the equivalent one-dimensional American put option has a risk-free rate of 0.06, a dividend yield of 0.02932, and a volatility of 0.14615. We set the strike price to $\Gamma=110$. The theoretical price of the corresponding European option is 9.649 according to the Black-Scholes formula, and the benchmark value for the one-dimensional American put option, calculated using the binomial tree method on Smirnov’s website with 20000 nodes, is 10.816. Table 3 presents the learning results obtained by our ML algorithm for different dimensions, which again demonstrate the high accuracy and scalability of the algorithm.
Dimension $d$ Last learned price by our method Runtime in sec.
40 10.828 1916
80 10.803 2948
120 10.813 4141
160 10.809 5665
200 10.795 7732
Table 3: Comparison of learned prices and runtime for different values of dimension $d$. The initial stock price $x_{0}=100^{1/\sqrt{d}}$, the time horizon $T=1$, the strike price $\Gamma=110$, the risk-free rate $\rho=0.06$, the volatility $\sigma_{i}=\frac{1}{\sqrt{10}}\min\{0.04[(i-1)\bmod 40],\,1.6-0.04[(i-1)\bmod 40]$ and dividend yield $\delta_{i}=\rho-\frac{\sum_{i=1}^{d}\sigma_{i}^{2}}{d^{2}}(i-\frac{1}{2})-\frac{1}{50\sqrt{d}}$ for each $i\in\left\{1,2,\ldots,d\right\}$. The payoff function when exercising the option is $g(x)=\left(\Gamma-\left[\prod_{k=1}^{d}|x^{(k)}|^{1/\sqrt{d}}\right]\right)^{+}$. The benchmark value for the price of this option is 10.816.
5.2 Optimally stopping a fractional Brownian motion
This example is adapted from Section 4.3 in Becker et al., (2019) and Section 4.2.3 in Dong, (2024), and is an optimal stopping problem for a fractional Brownian motion with Hurst parameter $H\in(0,1]$. By augmenting the state variable to include the entire history of the process, the problem becomes Markovian, and numerically a high-dimensional optimal stopping problem.
Specifically, a fractional Brownian motion with Hurst parameter $H\in(0,1]$ is a continuous, centered Gaussian process $(W_{t}^{H})_{t\geqslant 0}$ with covariance structure
$\mathbb{E}[W_{t}^{H}W_{s}^{H}]=\frac{1}{2}\left(t^{2H}+s^{2H}-|t-s|^{2H}\right),~~t,s\geqslant 0.$
When $H=1/2$, $W^{H}$ is a standard Brownian motion, and when $H\neq 1/2$, $W^{H}$ is neither a martingale nor a Markov process. The objective is to
$\max_{0\leqslant\tau\leqslant 1}\mathbb{E}[W_{\tau}^{H}]$
over all $W^{H}$-stopping times $\tau\in[0,1]$. This problem is non-Markovian and therefore falls outside the primary scope of this paper. However, we can approximate the optimal stopping value by discretizing the process. When combined with state space augmentation, this discretization transforms the problem into a Markovian one, thereby enabling the application of the martingale-based approach developed in this work.
To do this, denote $t_{n}=n/100$, for $n=0,1,\ldots,100$. For details on simulating trajectories of fractional Brownian motion, see Becker et al., (2019). Define the following 101-dimensional state process:
$X_{0}$ $=\left(W_{t_{0}}^{H},W_{t_{0}}^{H},W_{t_{0}}^{H},\ldots,W_{t_{0}}^{H}\right),\quad\quad\quad\quad\ X_{1}=\left(W_{t_{1}}^{H},W_{t_{0}}^{H},W_{t_{0}}^{H},\ldots,W_{t_{0}}^{H}\right),$
$X_{2}$ $=\left(W_{t_{2}}^{H},W_{t_{1}}^{H},W_{t_{0}}^{H},\ldots,W_{t_{0}}^{H}\right),\quad\ldots,\quad X_{100}=\left(W_{t_{100}}^{H},W_{t_{99}}^{H},W_{t_{98}}^{H},\ldots,W_{t_{0}}^{H}\right),$
and consider the following optimal stopping problem:
$\max_{\tau\in\mathcal{T}}\mathbb{E}[g(X_{\tau})],$
where $g(x)=x^{(1)}$ for a multi-dimensional vector $x$ with first component $x^{(1)}$, and $\mathcal{T}$ denotes the set of all $X$-stopping times. The multi-dimensional state process $X_{0}$, $X_{1}$, $\cdots$, $X_{100}$ is now Markovian, and the martingale-based approach applies.
In our simulation, we set the penalty factor $K$ to be 100 (corresponding to a time mesh grid $\Delta t=1/100$ as specified in Section 4.5.1), the temperature parameter $\lambda$ to be 0.1, the training batch size to be $2^{10}$, the testing batch size to be $2^{15}$, and the initial learning rate to 0.01 which is halved every 200 training steps. We train for 3000 steps using Algorithm 1.
Table 4 presents the learning results for different Hurst parameter values. The learned value is evaluated every 10 training steps, resulting in 300 observations for each case. The final observation is reported in the second column of Table 4. Note that in this example, if the agent does not stop until the terminal time, the stopping value is 0 in theory. Thus, our RL algorithm can achieve a high stopping value. Our learning results are comparable to benchmarks in the literature (Becker et al.,, 2019; Dong,, 2024), demonstrating high accuracy in learning the stopping decisions in a non-Markovian (or equivalently a high-dimensional) setting.
$H$ Last learned value by our method In Dong, (2024) In Becker et al., (2019) Runtime in sec.
0.01 1.496 1.513 1.519 3013
0.05 1.270 1.283 1.293 3008
0.10 1.027 1.043 1.049 3039
0.15 0.821 0.825 0.839 3476
0.20 0.641 0.651 0.658 3539
0.25 0.490 0.494 0.503 3519
0.30 0.360 0.364 0.370 3525
0.35 0.246 0.243 0.255 3532
0.40 0.148 0.144 0.156 3519
0.45 0.059 0.061 0.071 3519
0.50 0.002 0.000 0.002 3526
0.55 0.056 0.052 0.061 3528
0.60 0.110 0.108 0.117 3516
0.65 0.158 0.155 0.164 3528
0.70 0.201 0.192 0.207 3352
0.75 0.240 0.233 0.244 3080
0.80 0.274 0.261 0.277 3009
0.85 0.306 0.299 0.308 2454
0.90 0.337 0.331 0.337 2536
0.95 0.366 0.354 0.366 2370
Table 4: Comparison of stopping values and runtime for different values of Hurst parameter $H$ in the example of optimally stopping a fractional Brownian motion. The horizon length $T=1$.